OpenStudy (anonymous):

how many hours is 174 minutes?

OpenStudy (anonymous):

Divide 174 by 60.

OpenStudy (anonymous):

2.9 hours

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i think i should get the metal because i did the work

OpenStudy (zarkon):

\[174_\text{min}=174_\text{min}\frac{1_\text{hour}}{60_\text{min}}\] \[=174_\cancel{\text{min}}\frac{1_\text{hour}}{60_\cancel{\text{min}}}=\frac{174}{60}\text{hours}\]